# acmart.bib

### v2

@@ 3. Simulation
In this setting, there are 400 observations and up to seven distinct potential Sybil networks with probabilities between 0.1 and 0.9 of being Sybil and sizes between 10 and 90 observations. I simulate the setting 1000 times, and compare across four methods for generating $\hat{\beta}$, scoring by the mean-squared error between $\hat{\beta}$ and the known $\beta$. The four methods are: an estimator that includes all observations (“inclusion”), an estimator that excludes all suspected Sybil networks (“exclusion”), an estimator that excludes suspected networks with probabilities greater than 50% of being Sybil (“threshold”), and the optimal weighted estimator from Equation 2 (“weighted”).
Figure 1 shows the results. The exclusion-based estimator performs poorly, given how many observations it excludes. The inclusion-based and threshold-based estimators perform comparably to each other, trading off sample size with correlation in errors. However, the weighted estimator has markedly better performance in mean-squared error from the true coefficient value. This simple check demonstrates the potential value of the approach. By weighting potentially Sybil networks correctly and neither fully including nor excluding them, the estimator is more efficient.

### v3

@@ 3. Simulation
In this setting, there are 400 observations and up to seven distinct potential Sybil networks with probabilities between 0.1 and 0.9 of being Sybil and sizes between 10 and 90 observations. I simulate the setting 1000 times, and compare across six methods for generating $\hat{\beta}$, scoring by the mean-squared error between $\hat{\beta}$ and the known $\beta$. The six methods are: an estimator that excludes all suspected Sybil networks (“exclusion”), an estimator that includes all observations (“inclusion”), an estimator that excludes suspected networks with probabilities greater than 50% of being Sybil (“threshold”), an estimator that repeatedly re-samples entire networks by their probabilities of not being Sybil (“network-sampled”), an estimator that samples individual observations by their probabilities of not being Sybil (“observation-sampled”), and the optimal weighted estimator from Equation 2 (“weighted”).
Figure 1 shows the results. The exclusion-based estimator performs poorly, given how many observations it excludes. The other four benchmark estimators perform comparably, choosing different approaches to the trade-off between sample size and error correlation. However, the weighted estimator has markedly better performance in mean-squared error from the true coefficient value, compared to all other estimators. This is potentially especially surprising when compared to the estimators that sample networks and observations. However, the weight matrix is not a simple diagonal matrix that downweights suspected members of Sybil networks by probabilities and sizes; and it has off-diagonal elements that further adjust the estimates. Those are critically important, empirically.
This simple check demonstrates the potential value of the approach. By weighting potentially Sybil networks correctly and neither fully including nor excluding them, the estimator is more efficient.